OpenStudy (anonymous):

Find to consecutive integers whose product is 168.

OpenStudy (anonymous):

I just do not understand where to go from this equation: (x)(x+1)= 168, x^2 + x = 168

OpenStudy (anonymous):

x^2+x=168 i think you should use the quadratic formula to solve this

OpenStudy (anonymous):

how would the quadratic formula work if they have to be consecutive?

OpenStudy (anonymous):

2n(2n+1)=168 n(2n+1)=84 2n^2+n-84=0 now it become a quadratic equation solve it

OpenStudy (anonymous):

OKay sorry I'm just really confused... what is 2n?

OpenStudy (anonymous):

nothing...... i take here 2n and 2n+1 if u know apply here shri dhracharya formula

OpenStudy (anonymous):

shri dhracharya formula?

OpenStudy (raden):

the sum of a odd number and a even number impossible to get a even number, re-chek ur question...

OpenStudy (cwrw238):

x^2 + x - 168 = 0 can be used and the roots are not integers so there are no integers which satisfy this problem

OpenStudy (cwrw238):

oh hold on - its a product not a sum RadEn

OpenStudy (raden):

oh, sorry... i was mistaken

OpenStudy (cwrw238):

no worries

OpenStudy (cwrw238):

i think theres a mistake in question

OpenStudy (raden):

hmmm.. yea, i think so that

OpenStudy (cwrw238):

sure its not 156 ? instead of 168

OpenStudy (cwrw238):

i said that because it would give a result of 12 and 13